OpenStudy (anonymous):

3x over x plus 2 minus x over 5 Combine into one fraction and simplify

OpenStudy (anonymous):

\[\frac{3x}{x+2}-\frac{x}{5}\] \[=\frac{15x-x^2-2x}{5*(x+2)}\] \[=\frac{13x-x^2}{5(x+2)}\] \[=\frac{x(13-x)}{5(x+2)}\]